# Quantifying and Enhancing Multi-modal Robustness with Modality Preference

### v1

@@ 4.3 Validation for Effectiveness and scalability
We further provide experiment results of our methods extended to the Multi-Modal Transformer-based framework with hierarchical attention (Xu et al., 2023) (MMT) on the VGGS dataset. Both visual and audio class tokens are concatenated and linearly projected into the output space. We mainly evaluate the robustness under the Transformer model accurately, hence we train from scratch instead of using the pre-trained model. To validate our method, we only introduce the CRMT procedure step-1 on this transformer architecture as a comparison. That is because both audio and visual input influence each token, which is beyond our assumption. Based on the results in Table 3, our method can also improve the robustness of transformer-based architecture in most cases, which proves the broad prospects of our approach.
@@ 5 Conclusion
In this study, we present the essential components for multi-modal robustness and delve into the limitations imposed by modality preference. Additionally, we explore why multi-modal models exhibit vulnerabilities to attack specific modalities. Further, we introduce the Credible Multi-modal Robust Training procedure, a novel approach explicitly designed to boost the certified robustness.

### v2

@@ 4.3 Validation for Effectiveness and scalability
We further provide experiment results of our methods extended to the Multi-Modal Transformer-based framework with hierarchical attention (Xu et al., 2023a) (MMT) on the VGGS dataset. Both visual and audio class tokens are concatenated and linearly projected into the output space. We mainly evaluate the robustness under the Transformer model accurately, hence we train from scratch instead of using the pre-trained model. To validate our method, we only introduce the CRMT procedure step-1 on this transformer architecture as a comparison. That is because both audio and visual input influence each token, which is beyond our assumption. Based on the results in Table 3, our method can also improve the robustness of transformer-based architecture in most cases, which proves the broad prospects of our approach.
@@ 5 Conclusion
In this study, we present the essential components for multi-modal robustness and delve into the limitations imposed by modality preference. Additionally, we explore why multi-modal models exhibit vulnerabilities to attack specific modalities. Further, we introduce the Certifiable Robust Multi-modal Training procedure, a novel approach explicitly designed to boost the certified robustness.